# PerturBench: Benchmarking Machine Learning Models for Cellular Perturbation Analysis

### v1

@@ 5.2 Predicting Combinatorial Gene Overexpression Effects
The linear model performs relatively well in predicting combinatorial perturbationeffects, suggesting that many effects are approximately linearly additive in this dataset. The latent additive and decoder models were able to outperform the linear model in all metrics, suggesting they learned some non-linear interactions. The performance of the latent additive and decoder models are very similar, with the latent additive performing slightly better on rank metrics. This similarity suggests that the input gene expression values are not critical for strong model performance. CPA performs worse than the linear model in all metrics. SAMS-VAE underperforms with respect to all metrics, especially in the cosine LogFC metric. Biolord underperforms in the traditional cosine and RMSE metrics, but performs well with regards to the ranking metrics.
@@
SAMS-VAE $0.17\pm 2\times 10^{-3}$ $0.13\pm 5\times 10^{-4}$ $0.28\pm 1\times 10^{-2}$ $0.21\pm 1\times 10^{-2}$
@@ 5.3 Effect of Data Scaling
In this section we report the results of the data scaling scenario by verifying if models can take advantage of additional training data to better generalize perturbation effects across biological states. We base the analysis on the McFalineFigueroa23 dataset that includes both chemical and genetic perturbations across three cell lines. Since there were only 5 chemical perturbations, we consider each unique cell line and chemical perturbation a separate biological state, resulting in 15 total states with 525 genetic perturbations. To test whether adding biological states improves performance, we construct nested subsets of the dataset ($small\subset medium\subset full$), all sharing the same validation and test sets. Each of the subsets contains more biological states (details in Appendix E.4).
We find that model performance tends to improve with more training data in both cosine similarity and cosine rank (see Figure 2). A notable exception is our implementation of CPA, which does not improve on the cosine rank. This model also has the highest variance. The latent additive model seems to have the most favourable balance of performance on both cosine similarity and rank. Surprisingly, the linear model performs best on the rank metric. For further details please refer to Appendix D.2.
@@ 5.4 Effect of Data Imbalance
To test how different models’ performance is affected by data imbalance we downsample perturbations per cell type from Srivatsan20 to construct three sub-datasets with different levels of imbalance (Appendix E.4). The results are summarized in Figure 3. We observe that when the data is highly balanced, the linear model performs acceptably well, but this does not hold as imbalance increases. Imbalance may therefore be an important criteria for deciding the suitability of a linear model. CPA both with and without scGPT embeddings, is more robust to changes in data balance than the the Linear or latent additive models, however even with fully balanced data, the cosine rank is high indicating collapse. Interestingly, whilst the latent additive model is more markedly affected by data imbalance than other models, using scGPT embeddings seems to confer some buffer against this. This could be a highly useful feature when working with imbalanced perturb-seq data. The extent to which performance is affected by data imbalance highlights the importance of curated datasets and oversampling strategies (Cui et al.,, 2019).
@@ 6 Discussion
Further, we demonstrate that simple decoder-only models are able to explain much of the performance of perturbation models. This is surprising since those models only have access to one-hot encoded perturbation and covariate information. The, in principle, richer information present in the transriptomic readout hence has a smaller influence than the architecture and design choices of most models suggests (e.g. CPA and SAMS-VAE are designed around the expression data). Possible reasons include the lack of strong inductive biases (i.e. all neural networks are MLPs) or the noise inherent in single-cell RNA readouts.
There is a wide range of performance across the different tasks and datasets. This could be due to both the intrinsic difficulty of the task, and the amount of measurement or biological noise present in a dataset. For example, most models, including the linear model, perform well on the Norman19 combination prediction task (see Table 3), which suggests that this task is relatively easy as most combinations are linearly additive. Most models perform worse on the McFalineFigueroa23 data scaling and imbalanced data tasks, especially on the rank metrics (see Figures 2 and 3). For the data scaling task, most models performed better with more data, potentially suggesting perturbation models follow scaling laws (Kaplan et al.,, 2020). Thus, the data scaling and imbalanced data tasks may benefit from better modeling approaches, and we hope that these tasks will attract attention from the community.
@@
Given the diversity of dataloaders and model frameworks among public models, we aimed to assess the core components of each model. Thus, our benchmarking results should be interpreted as an assessment of how these core components perform rather than a perfectly accurate recreation of the public model implementation (see Appendix E.5 for details).
@@
In our codebase, we provide three main components: datasets and dataloaders, a model development framework, and an evaluation API with metrics (Appendix B). Each component can be used together or individually. For example, a user who wants to benchmark predictions generated by an existing model can use the evaluation API alone. Whereas a different user who wants a develop a new model with our model framework can use the entire codebase. Each component is also extensible, making it easy for users to add datasets, models, and metrics.
@@ Conclusion
The perturbation response modeling field holds great promise in complementing experimental approaches to identify novel disease targets and potential therapeutics. In this work we bring together some of the state-of-the-art models in a unified framework with thorough evaluation. We identify three key considerations in regards to benchmarking these models: (1) model components (which we tested through ablation studies), (2) performance on different datasets (especially at different scales and levels of imbalance), and (3) benchmarking metrics that fully characterize the effects of perturbations. We demonstrate the necessity of rank metrics and propose they become the standard in the field. Finally, we anticipate that our modular codebase will prove valuable in future model development and bencharmking efforts, ensuring meaningful discovery of potential targets to treat diseases.
@@
Appendix A Benchmarking Pipeline
In this section, we provide a high level overview of our perturbation modelling framework. An overview of the elements our contribution provides is given in Figure A.1. PerturBench includes A) processed datasets, B) functions for data processing and (meaningful) data splits, C) dataloaders for batched computation and model training, D) a variety of models and modeling elements and, finally E) an evaluation process.
Figure A.1: A) Single cell perturbational datasets at multiple scales. B) Biologically relevant covariate transfer and combinatorial prediction data splits. C) Dataloaders support two training modes: 1) supervised mode which involves mapping a sampled control cell to a perturbed cell and 2) unsupervised (or autoencoder) mode which maps a perturbed cell to itself. D) A model zoo with modular components such as relevant baseline models, adversarial loss, perturbation sparsity, and others. E) Standardized benchmarking suite supporting flexible pipelines and metrics for evaluating models
Appendix B Software Framework: The PerturBench Library
@@ B.3 Data Splitting
B.4 Model Abstraction: Model Base Class
@@ B.5 Evaluation
Appendix C Additional Modeling Background
@@
This can include, for instance, ensuring that the control cell is from the same cell type, experiment or batch. More complex methods involve using optimal transport to identify the control cell most likely to transition into a given perturbed cell (Jiang et al.,, 2024) or even to use optimal transport to assist in perturbation prediction (Bunne et al.,, 2023).
@@ C.1 Perturbation embeddings
Appendix D Further Results
D.1 Collapse and Rank Metrics
@@
Figure D.1: Cosine similarity matrix based on log-fold changes induced by perturbations A) Hyperparameter optimised DecoderOnly model with only cell type as covariates. B) Hyperparameter optimised DecoderOnly model with covariates and perturbations as input. C) Hyperparameter optimised version of our CPA fork. D) true log-fold changes in the underlying dataset. Clusters correspond to cell types A549, K562, MCF-7.
To further demonstrate this problem and the need for the rank metrics introduced in this paper, we will consider a simple example based on the Srivatsan20 dataset. This dataset contains the outcome of small molecules applied to three cancer cell lines: A549, K562, MCF-7. To simulate the impact mode collapse, we will investigate three models: 1) our DecoderOnly taking as input only the cell type (as a one-hot encoded vector), 2) our DecoderOnly taking as input covariates and perturbations, and 3) CPA taking as input covariates, perturbation and expression of perturbed cell. All models have been hyperparameter optimized with optuna, using 60 iterations. In addition, we will compare the model prediction to the data derived from the experimentally observed expression values. The result is visualized in Figure D.1 using matrix heatmaps and numerical results are presented in Table 4. Higher values (darker color) in the heatmaps correspond to higher similarity between predicted log-fold changes of different perturbation and cell type combinations on the validation data. Plot A) shows the DecoderOnly model based on the cell type alone. The large squares correspond to the three cell types and unsurprisingly, the prediction is identical for all perturbation based on the same cell type (as this information is not used by the model.) The true data, D), in contrast shows clusters, some of which are related to cell type, but overall different perturbations have different impacts. However, the model visualized in A) provides a valuable baseline, because it demonstrates what performance can be achieved without being able to provide any insight into the underlying perturbation modelling task. Notably, the RMSE are of roughly similar order of magnitude for all models, making it challenging to discern which level of performance is adequate, that is, there is no intrinsic quality in the value 0.027 in the DecoderOnly model that would suggest that the model is not capabale to predict perturbation effect at all.
@@
D.2 More scaling results
Most models have similar performance on the validation and test splits. However, CPA notably overfits to the validation split on the full data split and the linear model actually performs better on the test split than the validation split (see Figure D.2).
Figure D.2: Scaling of cosine similarity (left) and its rank (right) with increasing size of data included in the training process ($x$-axis) for several perturbation response models. Points represent results on test data for 5 different seeds, the line represent their average. Dotted lines are results on validation, solid lines are results on test.
Appendix E Implementation Details
E.1 Dataset Summary
Dataset 1 (Norman19).
@@ E.1 Dataset Summary
Dataset 3 (McFalineFigueroa23).
This dataset (McFaline-Figueroa et al.,, 2024) includes 525 genetic perturbations across 3 cell lines and 5 chemical treatments (which can be seen as 15 unique biological states). We selected this dataset due to the large number of perturbations and the fact that it contains multiple covariates (cell lines and chemical treatments). This dataset has not been previously used to benchmark perturbation prediction models and we believe that it adds a more realstic evaluation of the biological complexities that models will face in real-life deployment settings.
E.2 Dataset Curation
@@ E.5 Models
E.6 Hyperparameter Optimization
E.6.1 Identifying a Hyperparameter Metric
@@
  • Considering the results of Section D.1 hyperparameter optimization can not simply be carried out on one metric, such as RMSE, as we have established that this metric alone does not cover all aspects of model performance.
@@
E.6.2 Hyperparameter Ranges
For hyperparameter optimization we used optuna (Akiba et al.,, 2019). Hence, we can define all hyperparameter ranges as optuna distributions, either in the form of categorical, int or float. We describe the seed and the specific optuna hyperparameter ranges as well as their distribution classes in Tables 6, 7, 8, 9, 10 and 11.
@@ E.6.2 Hyperparameter Ranges
E.7 Compute Resources

### v2

@@ 5.3 Predicting Combinatorial Gene Overexpression Effects
The linear model performs relatively well in predicting combinatorial perturbation effects, suggesting that most perturbation effects are approximately linearly additive in this dataset. The latent additive and decoder models were able to outperform the linear model in all metrics, suggesting they learned some non-linear interactions. The performance of the latent additive and decoder models are very similar, with the latent additive performing slightly better on rank metrics. This similarity suggests that the input gene expression values are not critical for strong model performance. CPA and GEARS performs worse than the linear model in all metrics. Biolord and SAMS-VAE are worse than the linear model in the traditional cosine and RMSE metrics, but performs similarly with regards to the ranking metrics.
@@
SAMS-VAE $0.45\pm 2\times 10^{-2}$ $0.084\pm 7\times 10^{-4}$ $0.021\pm 5\times 10^{-3}$ $0.026\pm 2\times 10^{-3}$
@@ 5.4 Effect of Data Scaling
In this section we report the results of the data scaling scenario by verifying if models can take advantage of additional training data to better generalize perturbation effects across biological states. We base the analysis on the McFalineFigueroa23 dataset that includes both chemical and genetic perturbations across three cell lines. Since there were only 5 chemical perturbations, we consider each unique cell line and chemical perturbation a separate biological state, resulting in 15 total states with 525 genetic perturbations. To test whether adding biological states improves performance, we construct nested subsets of the dataset ($small\subset medium\subset full$), all sharing the same validation and test sets. Each of the subsets contains more biological states (details in Appendix D.4).
We find that model performance tends to improve with more training data in both cosine similarity and cosine rank (see Figure 3). A notable exception is our implementation of CPA, which does not improve on the cosine rank. This model also has the highest variance. The latent additive model seems to have the most favourable balance of performance on both cosine similarity and rank. Surprisingly, the linear model performs best on the rank metric. For further details please refer to Appendix C.2.
We further assessed how these models perform on large datasets with complex covariates by applying our benchmarking suite to the Jiang24 dataset, a large, 1.6 million cell dataset with complex covariates. The dataset contained 6 cell lines with 5 unique cytokine treatments, which we modeled as 30 distinct biological states and 219 genetic perturbations. However the set of perturbations applied was different for each cytokine treatment. We used a similar splitting strategy where we held out 70% of the covariates from 9 cell states for validation/testing.
The results are summarized in Table 6 where we see that the decoder model performs best for both the rank and RMSE/cosine metrics. CPA outperforms the Latent Additive model on the RMSE/cosine metrics and slightly underperforms on the rank metrics. The linear model is much worse than CPA and the latent additive model on the RMSE/cosine metrics, performs similarly on the RMSE rank metric, and is close to the decoder model on the cosine rank metric.
@@ 5.5 Effect of Data Imbalance
To test how different models’ performance is affected by data imbalance we downsample perturbations per cell type from Srivatsan20 to construct three sub-datasets with different levels of imbalance (Appendix D.4). The results are summarized in Figure 4. We observe that when the data is highly balanced, the linear model performs acceptably well, but this does not hold as imbalance increases. Imbalance may therefore be an important criteria for deciding the suitability of a linear model. CPA both with and without scGPT embeddings, is more robust to changes in data balance than the the Linear or latent additive models, however even with fully balanced data, the cosine rank is high indicating collapse. Interestingly, whilst the latent additive model is more markedly affected by data imbalance than other models, using scGPT embeddings seems to confer some buffer against this. This could be a highly useful feature when working with imbalanced perturb-seq data. The extent to which performance is affected by data imbalance highlights the importance of curated datasets and oversampling strategies (Cui et al.,, 2019).
@@ 6 Discussion
Further, we demonstrate that simple decoder-only models are able to explain much of the performance of perturbation models and for some datasets, are the best performing model. This is surprising since those models only have access to one-hot encoded perturbation and covariate information. The information present in the transriptomic readout hence has a smaller influence than the architecture and design choices of most models suggests (e.g. CPA and SAMS-VAE are designed around the expression data). Possible reasons include the lack of strong inductive biases (i.e. all neural networks are MLPs) and the noise inherent in either the single-cell RNA readouts or the perturbation labels. Specifically, CRISPR knockdown or knockout can be inefficient, resulting in some cells receiving the correct guide RNA but not having a robust gene knockdown or knockout (Liu et al.,, 2020). Additionally, high levels of heterogeneity in the control cells not captured by covariate labels could hinder the performance of the matching methods, as they randomly sample control cells to match to a given perturbed cell.
There is a wide range of performance across the different tasks and datasets. This could be due to both the intrinsic difficulty of the task, and the amount of measurement or biological noise present in a dataset. For example, the linear model performs well on the Norman19 combination prediction task (see Table 4), which suggests that this task is relatively easy as most combinations are linearly additive. Most models perform worse on the McFalineFigueroa23, Jiang24, data scaling and imbalanced data tasks, especially on the rank metrics (see Table 6, Figure 3, and Table 4). For the data scaling task, most models performed better with more data, potentially suggesting perturbation models follow scaling laws (Kaplan et al.,, 2020). Thus, the Jiang24, data scaling and imbalanced data tasks may benefit from better modeling approaches, and we hope that these tasks will attract attention from the community.
@@
Given the diversity of dataloaders and model frameworks among public models, we aimed to assess the core components of each model. Thus, our benchmarking results should be interpreted as an assessment of how these core components perform rather than a perfectly accurate recreation of the public model implementation (see Appendix D.5 for details).
@@
In our codebase, we provide three main components: datasets and dataloaders, a model development framework, and an evaluation API with metrics (Appendix A). Each component can be used together or individually. For example, a user who wants to benchmark predictions generated by an existing model can use the evaluation API alone. Whereas a different user who wants a develop a new model with our model framework can use the entire codebase. Each component is also extensible, making it easy for users to add datasets, models, and metrics.
@@ Conclusion
The perturbation response modeling field holds great promise in complementing experimental approaches to identify novel disease targets and potential therapeutics. In this work we bring together some of the state-of-the-art models in a unified framework with thorough evaluation. We identify three key considerations in regards to benchmarking these models: (1) model components (which we tested through ablation studies), (2) performance on different datasets (especially at different scales and levels of imbalance), and (3) benchmarking metrics that fully characterize the effects of perturbations. We demonstrate the necessity of rank metrics and propose they become the standard in the field. Finally, we anticipate that our modular codebase will prove valuable in future model development and benchmarking efforts, ensuring meaningful discovery of potential targets to treat diseases.
@@
Appendix A Software Framework: The PerturBench Library
@@ A.3 Data Splitting
A.4 Model Abstraction: Model Base Class
@@ A.5 Evaluation
Appendix B Additional Modeling Background
@@
This can include, for instance, ensuring that the control cell is from the same cell type, experiment or batch. More complex methods involve using optimal transport to identify the control cell most likely to transition into a given perturbed cell (Jiang et al., 2024b, ) or even to use optimal transport to assist in perturbation prediction (Bunne et al.,, 2023).
@@ B.1 Perturbation embeddings
Appendix C Further Results
C.1 Collapse and Rank Metrics
@@
Figure C.1: Cosine similarity matrix based on log-fold changes induced by perturbations A) Hyperparameter optimised DecoderOnly model with only cell type as covariates. B) Hyperparameter optimised DecoderOnly model with covariates and perturbations as input. C) Hyperparameter optimised version of our CPA fork. D) true log-fold changes in the underlying dataset. Clusters correspond to cell types A549, K562, MCF-7.
To further demonstrate this problem and the need for the rank metrics introduced in this paper, we will consider a simple example based on the Srivatsan20 dataset. This dataset contains the outcome of small molecules applied to three cancer cell lines: A549, K562, MCF-7. To simulate the impact mode collapse, we will investigate three models: 1) our DecoderOnly taking as input only the cell type (as a one-hot encoded vector), 2) our DecoderOnly taking as input covariates and perturbations, and 3) CPA taking as input covariates, perturbation and expression of perturbed cell. All models have been hyperparameter optimized with optuna, using 60 iterations. In addition, we will compare the model prediction to the data derived from the experimentally observed expression values. The result is visualized in Figure C.1 using matrix heatmaps and numerical results are presented in Table 5. Higher values (darker color) in the heatmaps correspond to higher similarity between predicted log-fold changes of different perturbation and cell type combinations on the validation data. Plot A) shows the DecoderOnly model based on the cell type alone. The large squares correspond to the three cell types and unsurprisingly, the prediction is identical for all perturbation based on the same cell type (as this information is not used by the model.) The true data, D), in contrast shows clusters, some of which are related to cell type, but overall different perturbations have different impacts. However, the model visualized in A) provides a valuable baseline, because it demonstrates what performance can be achieved without being able to provide any insight into the underlying perturbation modelling task. Notably, the RMSE are of roughly similar order of magnitude for all models, making it challenging to discern which level of performance is adequate, that is, there is no intrinsic quality in the value 0.027 in the DecoderOnly model that would suggest that the model is not capabale to predict perturbation effect at all.
@@
C.2 More scaling results
Most models have similar performance on the validation and test splits. However, CPA notably overfits to the validation split on the full data split and the linear model actually performs better on the test split than the validation split (see Figure C.2).
Figure C.2: Scaling of cosine similarity (left) and its rank (right) with increasing size of data included in the training process ($x$-axis) for several perturbation response models. Points represent results on test data for 5 different seeds, the line represent their average. Dotted lines are results on validation, solid lines are results on test.
C.3 Additional results on a large perturbation response dataset
Table 6: Results of the Jiang24 experiment which contains 30 different cell states (6 cell lines & 5 cytokine treatments), with 70% of perturbations from 9 cell states held out for validation/testing. Results are reported as mean $\pm$ one standard deviation. Best performance per metric is indicated in bold.
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.58\pm 4\times 10^{-2}$ $0.017\pm 5\times 10^{-4}$ $0.40\pm 8\times 10^{-3}$ $0.42\pm 9\times 10^{-3}$
LA $0.47\pm 1\times 10^{-3}$ $\mathbf{0.015\pm 7\times 10^{-5}}$ $0.38\pm 6\times 10^{-3}$ $0.38\pm 7\times 10^{-3}$
Decoder $\mathbf{0.64\pm 8\times 10^{-4}}$ $\mathbf{0.015\pm 3\times 10^{-5}}$ $\mathbf{0.32\pm 8\times 10^{-3}}$ $\mathbf{0.32\pm 5\times 10^{-3}}$
Linear $0.17\pm 8\times 10^{-5}$ $0.038\pm 5\times 10^{-5}$ $0.34\pm 2\times 10^{-3}$ $0.43\pm 1\times 10^{-3}$
Appendix D Implementation Details
D.1 Dataset Summary
Dataset 1 (Norman19)
@@ D.1 Dataset Summary
Dataset 3 (Frangieh21)
[…]
Dataset 4 (Jiang24)
This dataset (Jiang et al., 2024a, ) includes 219 genetic perturbations across 6 cell lines and 5 cytokine treatments (which can be seen as 30 unique biological states). We selected this dataset due to the large number of biological states and the fact that the perturbations were chosen because they had been reported to modulate cytokine signaling. This dataset has not been previously used to benchmark perturbation prediction models.
Dataset 5 (McFalineFigueroa23)
This dataset (McFaline-Figueroa et al.,, 2024) includes 525 genetic perturbations across 3 cell lines and 5 chemical treatments (which can be seen as 15 unique biological states). We selected this dataset due to the large number of perturbations and the fact that it contains multiple covariates (cell lines and chemical treatments). This dataset has not been previously used to benchmark perturbation prediction models.
D.2 Dataset Curation
@@ D.4 Data Splitting
Jiang24 splits
We held out 70% of perturbations in all 12 combinations of the following cytokines: IFNG, INS, TGFB and cell lines: k562, mcf7, ht29, hap1. The remaining perturbations were used for training.
The attached codebase has a python notebook responsible for generating this split: notebooks/neurips2024/build_jiang24_frangieh21_splits.ipynb.
Frangieh21 splits
We held out 70% of the perturbations in the co-culture condition and trained on the remaining perturbations.
The attached codebase has a python notebook responsible for generating this split: notebooks/neurips2024/build_jiang24_frangieh21_splits.ipynb.
@@ D.5 Models
GEARS
Since the GEARS model differs from the other models in its use of GNNs to encode gene expression values and perturbations and since the authors did not recommend applying it to the covariate transfer task, we chose not to reimplement GEARS in our repo. We instead wrote a helper function and HPO script for training and evaluating GEARS using the publicly available version on the same Norman19 split we used for the other models.
@@
D.6 Hyperparameter Optimization
D.6.1 Identifying a Hyperparameter Metric
@@
  • Considering the results of Section C.1 hyperparameter optimization can not simply be carried out on one metric, such as RMSE, as we have established that this metric alone does not cover all aspects of model performance.
@@
D.6.2 Hyperparameter Ranges
For hyperparameter optimization we used optuna (Akiba et al.,, 2019). Hence, we can define all hyperparameter ranges as optuna distributions, either in the form of categorical, int or float. We describe the seed and the specific optuna hyperparameter ranges as well as their distribution classes in Tables 8, 9, 10, 11, 12 and 14.
@@ D.6.2 Hyperparameter Ranges
Table 14: GEARS hyperparameter range.
Hyperparameter Distribution
Number of layers in perturbation GNN:
num_go_gnn_layers Int: 1 to 3, , step=1
Number of layers in gene GNN:
num_gene_gnn_layers Int: 1 to 3, step=1
Number of neighboring perturbations in GO graph:
num_similar_genes_go_graph Int: 10 to 30, step=10
Number of neighboring genes in gene co-expression graph:
num_similar_genes_co_express_graph Int: 10 to 30, step=10
Width of the encoder:
hidden_size Int: 32 to 512, step=32
Minimum coexpression threshold:
co_express_threshold_graph Float: 0.2 to 0.5, step=0.1
Learning rate:
lr Float: 5e-6 to 5e-3, log=True
Weight decay:
wd Float: 1e-8 to 1e-3, log=True
D.7 Compute Resources